Find the area of the largest isosceles triangle that can be inscribed in a circle of radius 6 inches.


1
Expert's answer
2022-01-11T18:46:14-0500


From the diagram, the height of the triangle is "6+h"

We can as well gfrac{det the base using Pythagoras theorem.

"b=2\sqrt{36-h^2}"

Thus, the area of the triangle is

"A=\frac{1}{2}(2\sqrt{36-h^2})(6+h)=(\sqrt{36-h^2}(6+h)"

Find the derivative of A and set it to 0.

"\frac{dA}{dh}=\sqrt{36-h^2}-\frac{h(6+h)}{\sqrt{36-h^2}}=0\\
36-h^2-h(6+h)=0\\
36-2h^2-6h=0\\
h^2+3h-18=0\\
(h-3)(h+6)=0\\
h=3,h=-6"

Thus h=-6,3 are the critical points.

"\frac{d^2A}{dh^2}" at h=-6 >0, "\frac{d^2A}{dh^2}" at h=3<0.

Thus, "h=3" is the maximum point.

The largest area of the triangle is

"A=(\sqrt{36-3^2})(6+3)=27\sqrt{3} in^2"